OpenStudy (anonymous):

Show the steps to your solution.Find the value of x in the equation 5^x=1/125

OpenStudy (anonymous):

5^x = 1 / 5^3 5^x * 5^3 = 1 5^x * 5^3 = 5^0 x + 3 = 0 x = -3

OpenStudy (anonymous):

\[x^{5} = {1 \over 125}\]\[x = \sqrt[5]{1 \over 125} = ({1 \over 125})^{1/5}\]\[x = {1 \over 5}\]

OpenStudy (anonymous):

I don't think you can write 5^x like x^5!

OpenStudy (anonymous):

oh my bad, misread it